# Some insights into depth estimators for location and scatter in the multivariate setting

### v1

@@ 3 Maximum bias and concentration inequalities
$\inf_{\boldsymbol{\theta},\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(P_{0}\right)}P\left(\left\|\boldsymbol{\hat{\theta}}_{T}-\boldsymbol{\theta}\right\|^{2}\leq C\left(\max\left\{\frac{p}{n},\varepsilon^{2}\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha$ (4)
(Theorem 2.1 of Chen et al. (2018)). The constant $C$ in the concentration inequality (4) is actually affected by the asymptotic maximum bias of Tukey’s median.
However, the bound (4) can be derived in a more illuminating manner by explicitly incorporating the maximum bias, as the maximum bias governs the behavior of the estimator when the sample size is sufficiently large. Without enlarging the bound in (4), we obtain a more informative inequality,
$\inf_{\boldsymbol{\theta},\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(P_{0}\right)}P\left(\left\|\hat{\boldsymbol{\theta}}_{T}-\boldsymbol{\theta}\right\|^{2}\leq\tilde{C}\left(\max\left\{\frac{p}{n},B^{2}\left(\mathbf{\hat{\theta}}_{T},\varepsilon,\Phi\right)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
(see Lemma 9 in the Appendix).
The reasoning behind expecting the asymptotic maximum bias function to appear in the concentration inequality is as follows: As the sample size tends to infinity, the estimator is expected to converge to the functional value $\hat{\boldsymbol{\theta}}_{T}\left(P\right).$The quantity $\left\|\hat{\boldsymbol{\theta}}_{T}\left(P\right)-\boldsymbol{\theta}\right\|$ remains within a range bounded above by the maximum bias corresponding to the given level of contamination, since the distributions vary over the entire $\varepsilon$-contamination neighborhood.
On the other hand, Chen et al. (2018) also derived a concentration inequality for the deepest estimator for the dispersion matrix. They showed that, with probability at least $1-2\delta,$
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(F_{0}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C\left(\max\left\{\frac{p}{n},\varepsilon^{2}\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$ (5)
With a similar reasoning to that of Lemma 9 we can obtain that (see Lemma 10 below)
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(F_{0}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C^{\ast}\left(\max\left\{\frac{p}{n},B_{E}^{2}(\varepsilon)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
with $B_{E}\left(\varepsilon\right)=\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right].$ By these means, the concentration inequality is also able to uncover the maximum bias of the deepest as it will be shown in the next Section.
4 Maximum bias for the deepest scatter matrix
From now onwards we will assume that
The bounds found using the concentration inequalities of Chen et al. (2018) suggest to study the asymptotic maximum bias of the deepest scatter matrix to verify the accuracy of the bounds. We will summaryze the steps to get the maximum bias function for the deepest scatter matrix. First of all we compute the depth of any matrix under the normal model and the deepest estimator under this distribution, which is given by $\Phi^{-1}\left(3/4\right)I.$ We next verify that in case of having a sequence of contaminations in the $\varepsilon-$contamination neighborhood which yields the deepest estimator to have either the largest eigenvalue going to infinity or the smallest one going to $0,$ then the level of contamination $\varepsilon\geq 1/3,$ which entails that $\varepsilon^{\ast}\geq 1/3.$ Next, we calculate the depth of any matrix in $\mathcal{E}$ under point mass contaminations located along the direction given by a vector $\mathbf{e\in\mathbb{R}}^{p}$. Moreover we consider the set of the matrices whose eigenvector associated with the largest eigenvalue coincides with the vector $\mathbf{e}$, which yields a more specific formula rather the general one for any matrix given in the previous step. Then, in this class, we calculate a deepest scatter matrix (since uniqueness cannot be suspected at all) in which the largest and smallest eigenvalues  attainable for point mass contaminations coincide with those of the bounds found in the concentration inequalities. On the one hand we show that in case of having a sequence of matrices such that the maximum eigenvalue tends to infinity its depth could only converge to a values less or equal than $\min\left(\varepsilon,1-\varepsilon\right).$ On the other hand, we prove that the deepest estimator in the case of using point mass contaminations is less or equal to $\left(1-\varepsilon\right)/2$. Therefore we can conclude that if we have a level of contamination $\varepsilon$ such that the deepest estimator have smallest and largest eigenvalues bounded above and below for any contamination, the level of contamination must be less or equal to $1/3$ and we get that the breakdown $\varepsilon^{\ast}=1/3,$ which coincide with that of Tukey’s median. The reminiscence of Tukey’s median is even emphasized since we can use a similar reasoning to that of Chen and Tyler (2002) to get a bound for the asymptotic maximum bias curve. And the bound found let us confirm that it agrees with that of the bounds obtained in the concentration inequalities given in the former Section, that is
Theorem 2
All the proofs are deferred to the Appendix.

### v2

@@ 6.1 The estimators
    $(\widehat{\boldsymbol{\mu}},\widetilde{\Sigma})=\arg\min_{u\in\mathbb{R}^{p},\Sigma\in\mathcal{E},|\Sigma|=1}\sum_{i=1}^{n}\rho\left(\frac{d_{i}}{cS^{0}}\right),$
[…]
    $\rho_{\mathrm{SHR}}(d)=\frac{1}{6.494}\begin{cases}d&\text{if }d\leq 4,\\ s(d)&\text{if }4<d\leq 9,\\ 6.494&\text{if }d>9,\end{cases}$
    where
    $s(d)=3.534-1.944\,d+0.864\,d^{2}-0.104\,d^{3}+0.004\,d^{4}.$
    For more details, see Maronna and Yohai (2017).
  • 7.
    The Stahel-Donoho location scatter estimator (SD) was proposed by Stahel (1981) and Donoho (1982), weighting the observations according to a notion of outlyingness. Given $\mathbf{u}\in\mathcal{S}^{p-1}$, denote by $\mathbf{u}^{\prime}\mathbb{X}=\left\{\mathbf{u}^{\prime}\mathbf{x}_{1},\ldots,\mathbf{u}^{\prime}\mathbf{x}_{n}\right\}$. The outlyingness with respect to $\mathbb{X}$ of $\mathbf{x}\in\mathbb{R}^{p}$ along $\mathbf{u}$ is defined by
    $t(\mathbf{x},\mathbf{u})=\frac{\mathbf{x}^{\prime}\mathbf{u}-L\left(\mathbf{u}^{\prime}X\right)}{S\left(\mathbf{u}^{\prime}X\right)},$
    where $L$ and $S$ are robust univariate location and scale estimators, for example the median and the MAD respectively. Now, the outlyingness of $\mathbf{x}$ is defined by
    $t(\mathbf{x})=\sup_{\mathbf{u}\in\mathcal{S}^{p-1}}|t(\mathbf{x},\mathbf{u})|.$
[…]
    $\widehat{\boldsymbol{\mu}}$ $=$ $\frac{1}{\sum_{i=1}^{n}w_{i1}}\sum_{i=1}^{n}w_{i1}\mathbf{x}_{i},$
    $\widehat{\Sigma}$ $=$ $\frac{1}{\sum_{i=1}^{n}w_{i2}}\sum_{i=1}^{n}w_{i2}\left(\mathbf{x}_{i}-\widehat{\boldsymbol{\mu}}\right)\left(\mathbf{x}_{i}-\widehat{\boldsymbol{\mu}}\right)^{t}$
    It is computed using the function CovSde from the rrcov package. This function uses the piecewise polynomial weight $W_{\mathrm{opt}}$ described in Maronna and Yohai (2017).
  • 8.
    The Deepest Estimator (MDepth) proposed by Chen, Gao and Ren (2018a). We use the function matrix_depth_by_descent from the R package DepthDescent. For more details see Chen, Gao and Ren (2018b).
We do not compute Tyler’s M-estimator of scatter, since we are only using routines available at the CRAN Project and the R package ICSNP takes the sample mean as multivariate location estimator by default, which makes the comparison with other robust proposals unreliable.
6.2 Contamination scenarios
Let $\mathcal{P}_{\varepsilon}\left(P_{0}\right)$ be the $\varepsilon$-contamination neighborhood with $P_{0}=N(\boldsymbol{\theta},\Sigma)$, $\boldsymbol{\theta}=\mathbf{0}$ and $\Sigma=I_{p}$, the $p\times p$ identity matrix. We assume that the contaminating distribution is $G_{k}=\delta_{x_{k}}$, a point mass at $\mathbf{Z}_{k}=(k,\ldots,k)^{t}$, with $k\in\mathbb{N}\cup\{0\}$. Consider $\boldsymbol{X}^{0}=(X_{1}^{0},\ldots,X_{p}^{0})^{\prime}\sim P_{0}$ and $B\sim Ber\left(\varepsilon\right).$ If $\boldsymbol{X}^{0}$ and $B$ are independent then
$\boldsymbol{X}=(1-B)\boldsymbol{X}^{0}+B\boldsymbol{Z}_{k}\sim P_{k},$ (20)
with $P_{k}=(1-\varepsilon)P_{\mathbf{0}}\left(\cdot\right)+\varepsilon G_{k}(\cdot)\in\mathcal{P}_{\varepsilon}\left(P_{0}\right).$ For a contamination rate $\varepsilon$ and a fixed constant $k$, using (20), we generate a sample $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ from the random vector $\mathbf{X}\sim P_{k}$.
We will consider the following scenarios:
  • Contamination proportions $\varepsilon=0.1$, $0.2$.
  • $k$ $\in$ $\left\{0,1,5,10,15,20,25\right\}$.
  • Dimensions $p=2,5$, 1010, 1515.
  • Sample size: $n=10p$ (see Maronna and Yohai (2017), Table 3), $n=40p$, $n=500p$.
For each combination of $p$, $\varepsilon$, and $k$, we generate $R=50$ independent data sets, denoted as $\mathbb{X}_{r}=\{\mathbf{x}_{1}^{(r)},\ldots,\mathbf{x}_{n}^{(r)}\}$, for $r=1,\ldots,R$.
6.3 Empirical bias
The effect of the distortion caused by having $P\in\mathcal{P}_{\varepsilon}\left(P_{0}\right)$ was measured through formulas (3.1) and (3.1). Following Maronna and Yohai (1995), we also use the condition number (CN) of an estimator $\hat{\Gamma}$ as an additional measure of bias, abbreviated as $B_{\text{CN}}$, and defined as the supremum of $\text{CN associated with }\hat{\Gamma}$ over the contamination neighbourhood.
To obtain empirical versions of $B$ and $B_{\text{CN}}$, we proceed as follows:
  • Step 1. For every $k=0,1,\ldots,25$, generate $R=50$ independent data sets, denoted by
    $\mathbb{X}_{k}^{(r)}=\{\mathbf{x}_{1}^{(r)},\ldots,\mathbf{x}_{n}^{(r)}\},\quad r=1,\ldots,R,$
    where each $\mathbb{X}_{k}^{(r)}$ is a sample from the distribution $P_{k}$. Let $\hat{\Gamma}_{k}^{(r)}$ denote the corresponding estimated scatter matrix. For every $\hat{\Gamma}_{k}^{(r)}$, compute $\hat{b}_{E,k}^{(r)}$ and $\hat{b}_{I,k}^{(r)}$, $r=1,\ldots,R$, by using the empirical version of $P_{k}$ corresponding to $\mathbb{X}_{k}^{(r)}$, denoted by $P_{n,k}^{(r)}$. Then, define the bias of the scatter matrix $\hat{\Gamma}_{k}^{(r)}$, based on the $r$-th sample from $P_{k}$, as
    $\hat{b}_{k}^{(r)}=\max\left\{\hat{b}_{I,k}^{(r)},\hat{b}_{E,k}^{(r)}\right\}=\max\left\{\lambda_{(1)}^{(r)},\lambda_{(p)}^{(r)-1}\right\},$
    where $\lambda_{(1)}^{(r)}$ and $\lambda_{(p)}^{(r)}$ denote the largest and the smallest eigenvalues of $\hat{\Gamma}_{k}^{(r)}$. For $1\leq r\leq R,$ let $_{k}^{(r)}=CN$$\left(\hat{\Gamma}_{k}^{(r)}\right)=\lambda_{(1)}^{(r)}/\lambda_{(p)}^{(r)}$ be the condition number of the estimated scatter matrix $\hat{\Gamma}_{k}^{(r)}.$ If $l$ is a location estimator, then $\hat{b}_{k}$ and CNk stand for
    $\hat{b}_{k}=l\left(\{\hat{b}_{k}^{(r)}\}\right),$ (21)
    $CN_{k}=l\{CN_{k}^{(r)}\}.$ (22)
    Step 2. The measures
    $\hat{B}=\max_{0\leq k\leq 25}\{\hat{b}_{i,k}\}\quad\text{and}\quad\hat{B}_{\text{CN}}=\max_{0\leq k\leq 25}\{CN_{k}\},$ (23)
    are empirical approximations to the maximum biases $B$ and $B_{\text{CN}}$, respectively.
In order to place the measures $\hat{b}_{I,k}^{(r)},\hat{b}_{E,k}^{(r)}$ and $CN_{k}^{(r)}$on a more comparable scale, we consider their logarithms. When sampling from a contaminated distribution $P_{k},$ the empirical distributions of $\log\left(\hat{b}_{I,k}^{(r)}\right),\log\left(\hat{b}_{E,k}^{(r)}\right)$ and $\log\left(CN_{k}^{(r)}\right)$ exhibit asymmetry and the presence of outliers, as shown by the boxplots in Figures 1 to 4. Therefore, it is more appropriate to use the median as the location measure $l$ in (21) and (22) rather than the mean. For the sake of simplicity, we refer to $\hat{B}$ and $\hat{B}_{\mathrm{CN}}$ as the “maximum medians of $B$ and CNCN”, respectively in the tables and figures. Further details are provided in the Section 6.5.
6.4 Efficiency
When the data come from the symmetric central model, the empirical distributions of $\log\left(CN_{k}^{(r)}\right)$ and $\log\left(b_{k}^{(r)}\right)$ tend to be more symmetric, and the mean is therefore more representative. Thus, given a scatter estimator, we can take $\hat{B}$ and $\hat{B}_{\mathrm{CN}}$ as defined in (23), but using the mean rather than the median in (21) and (22), which will be called the (empirical) mean absolute error, abbreviated as MAE. Furthermore, if $\hat{\Gamma}$ denotes a scatter estimator and $S_{n}$ denotes the sample covariance matrix, the ratio of the MAE for $S_{n}$ to the corresponding value for $\hat{\Gamma}$
[…]
may be considered a measure of efficiency at the normal central model $P_{0}$; see Maronna and Yohai (1995).
6.5 Results and conclusions
In the Tables 1 to 4 the efficiencies of scatter estimators are computed for different combinations of $n$ and $p$. For most estimators, efficiency remains stable or even slightly increases with growing $p$, suggesting favorable finite-sample behavior when the ratio $p/n$ is moderate. In particular, the MM estimator maintains high efficiency across all considered dimensions and sample sizes.
For the S-E, as can be seen in Tables 14, for small dimensions the efficiency is low, but it increases considerably as $p$ grows. This is a phenomenon described in the literature; see, for instance, Section 6.4.4. of Maronna et al. (2018). As discussed therein, when $p$ is large enough, almost all observations receive similar weights, except for observations far from the bulk of the data, yielding an estimator which closely approximates the sample covariance matrix. But increasing efficiency usually entails a loss of robustness when the dimension grows.
n MVE MCD SE ROCKE $M$ SD MDEPTH
$p=2$ 0.254 0.688 0.541 0.359 0.917 0.589 0.647
$p=5$ 0.339 0.736 0.826 0.565 0.941 0.601 0.798
$p=10$ 0.381 0.700 0.910 0.524 0.958 0.585 0.774
Table 1: Eff in log-scale based on CN (means) for $n=50$ over dimensions.
MVE MCD SE ROCKE $M$ SD MDEPTH
$p=2$ 0.446 0.690 0.580 0.433 0.806 0.617 0.282
$p=5$ 0.328 0.687 0.803 0.518 0.919 0.564 0.399
$p=10$ 0.355 0.678 0.876 0.444 0.952 0.582 0.471
Table 2: Eff in log-scale based on $B$ (means) for $n=50$ over dimensions.
n MVE MCD SE ROCKE $M$ SD MDEPTH
$p=2$ 0.224 0.630 0.515 0.378 0.944 0.567 0.630
$p=5$ 0.303 0.825 0.855 0.635 0.940 0.723 0.770
$p=10$ 0.385 0.856 0.951 0.706 0.951 0.786 0.846
$p=15$ 0.416 0.858 0.970 0.698 0.933 0.746 0.843
Table 3: Eff in log-scale based on CN (means) for $n=200$ over dimensions.
MVE MCD SE ROCKE $M$ SD MDEPTH
$p=2$ 0.368 0.667 0.583 0.452 0.875 0.636 0.151
$p=5$ 0.308 0.800 0.848 0.622 0.903 0.718 0.255
$p=10$ 0.339 0.811 0.935 0.652 0.935 0.754 0.331
$p=15$ 0.364 0.821 0.965 0.640 0.917 0.714 0.379
Table 4: Eff in log-scale based on $B$ (means) for $n=200$ over dimensions.
Figures 1 to 4 show boxplots of the biases $\{\hat{b}_{k}^{(r)}\}$, on a log scale, across the $R$ replicates for each scatter estimator, under selected scenarios (combinations of $p$, $n$, and $k$). The boxplots highlight the need for using medians rather than means as representative measures of the behavior of the bias measures. Note that most empirical distributions exhibit heavy tails, skewness, and outliers, especially when the contamination level is $\varepsilon=0.2$.
Figures 5 to 8 display the behavior of $\hat{b}_{k}=\text{median}_{1\leq r\leq R}\{\hat{b}_{k}^{(r)}\}$ (see (21) and (22)), for each scatter estimator, as functions of $k$, for selected dimensions $p$ and sample sizes $n$, under both contamination levels. Tables 5 to 12 show the maximum medians of $CN$ and $B$ for each scatter estimator across sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=2,\ldots,15$. From the tables we can conclude that MM tends to give the smallest maximum medians of $B$ for small and moderate sample sizes and for lower dimensions, showing the best overall performance. ROCKE often outperforms MM for larger sample sizes and in higher dimensions ($p\geq 10$). The figures suggest that, in low-dimensional settings ($p=2$), MM and MCD perform best. For larger sample sizes, MM remains the best-performing estimator, followed by SE and ROCKE, while MCD also shows competitive performance. In moderate and higher dimensions ($p=5,10$), MM consistently exhibits strong performance across contamination levels. Generally speaking, the bias curves decrease as $k$ grows, reflecting the robustness of these estimators to extreme outliers.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
20 0.10 4.72 2.51 1.13 1.58 1.93 0.69 1.24 0.97
20 0.20 5.36 3.03 1.95 2.35 2.38 0.81 1.69 1.32
80 0.10 4.81 1.49 0.60 0.79 0.94 0.40 0.72 0.62
80 0.20 5.54 1.71 1.26 1.29 1.40 0.46 1.14 1.05
1000 0.10 4.86 0.98 0.63 0.61 0.80 0.24 0.52 0.48
1000 0.20 5.55 1.35 1.33 1.17 1.30 0.45 1.02 0.96
Table 5: Empirical maximum bias $\hat{\text{B}}_{\text{CN}}$ (log scale) for each scatter estimator across sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=2$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
20 0.10 4.77 1.29 1.37 1.48 1.65 0.78 1.30 1.65
20 0.20 5.35 1.51 1.67 1.90 2.20 0.82 1.60 1.82
80 0.10 4.74 0.75 0.48 0.51 0.65 0.34 0.48 1.18
80 0.20 5.32 0.86 1.14 0.99 1.14 0.52 0.93 1.62
1000 0.10 4.76 0.49 0.48 0.41 0.51 0.21 0.40 1.14
1000 0.20 5.31 0.68 1.18 0.80 0.90 0.52 0.77 1.52
Table 6: Empirical maximum bias $\hat{B}$ (log scale) for each scatter estimator across sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=2$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
50 0.10 6.30 3.26 1.73 1.40 2.17 1.14 2.09 1.60
50 0.20 7.07 4.15 2.97 2.34 3.00 1.20 3.51 2.22
200 0.10 6.00 1.98 1.04 0.99 1.18 0.72 1.27 1.04
200 0.20 6.65 2.34 1.87 1.51 1.58 1.11 2.09 1.77
2500 0.10 5.81 1.57 0.66 0.76 0.75 0.52 0.82 0.92
2500 0.20 6.50 1.86 1.31 1.26 1.09 0.86 1.53 1.67
Table 7: Empirical maximum bias $\hat{\text{B}}_{\text{CN}}$ (log scale) for each scatter estimator under contamination levels $\varepsilon$ and different sample sizes $n$ for dimension $p=5$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
50 0.10 5.75 1.97 1.24 0.89 1.55 0.69 1.34 1.73
50 0.20 6.24 2.41 2.35 1.42 2.26 0.78 2.09 2.32
200 0.10 5.69 1.08 0.67 0.55 0.66 0.41 0.70 1.37
200 0.20 6.22 1.27 1.34 0.87 0.96 0.63 1.23 1.82
2500 0.10 5.65 0.84 0.36 0.49 0.46 0.29 0.53 1.15
2500 0.20 6.22 1.02 0.80 0.83 0.60 0.47 1.04 1.54
Table 8: Empirical maximum bias $\hat{B}$ (log scale) for each scatter estimator under contamination levels $\varepsilon$ and different sample sizes $n$ for dimension $p=5$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
100 0.10 7.00 3.32 2.06 1.64 1.92 1.40 2.52 2.52
100 0.20 7.80 4.56 3.35 2.24 2.48 1.58 4.41 3.97
400 0.10 6.70 2.23 1.29 1.30 1.14 1.09 1.51 2.03
400 0.20 7.41 2.75 2.33 1.86 1.39 1.56 2.64 3.59
5000 0.10 6.51 2.11 0.94 1.06 0.75 0.85 1.11 1.79
5000 0.20 7.22 2.48 1.79 1.64 0.97 1.34 2.06 3.34
Table 9: Empirical maximum bias $\hat{\text{B}}_{\text{CN}}$ (log scale) for each scatter estimator under different sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=10$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
100 0.10 6.34 1.87 1.40 0.90 1.33 0.77 1.35 2.04
100 0.20 6.88 2.41 2.49 1.31 1.79 0.81 2.67 2.39
400 0.10 6.30 1.26 0.69 0.86 0.61 0.66 0.97 1.53
400 0.20 6.89 1.48 1.50 1.31 0.77 1.01 1.85 2.32
5000 0.10 6.33 1.13 0.65 0.85 0.53 0.65 0.87 1.17
5000 0.20 6.91 1.40 1.01 1.34 0.62 1.04 1.69 2.00
Table 10: Empirical maximum bias $\hat{B}$ (log scale) for each scatter estimator under different sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=10$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
150 0.10 7.48 3.33 2.41 1.92 1.95 1.64 3.03 3.60
150 0.20 8.22 4.57 3.87 2.67 2.43 1.93 5.08 5.66
600 0.10 7.15 3.33 2.41 1.92 1.95 1.37 3.03 3.60
600 0.20 7.88 3.24 2.76 2.16 1.29 1.87 3.11 5.19
7500 0.10 6.92 2.46 1.17 1.30 0.69 1.10 1.06 2.25
7500 0.20 7.62 3.00 2.15 1.91 0.84 1.65 1.73 4.77
Table 11: Empirical maximum bias $\hat{\text{B}}_{\text{CN}}$ (log scale) for each scatter estimator under different sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=15$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
150 0.10 6.68 1.96 1.37 1.10 1.32 0.87 1.79 2.37
150 0.20 7.28 2.39 2.43 1.69 1.77 1.05 3.47 2.62
600 0.10 6.73 1.47 0.92 1.15 0.56 0.93 1.31 2.23
600 0.20 7.31 1.77 1.56 1.65 0.71 1.35 2.36 1.90
7500 0.10 6.73 1.33 0.88 1.12 0.50 0.92 0.89 1.23
7500 0.20 7.31 1.70 1.35 1.66 0.55 1.39 1.46 2.02
Table 12: Empirical maximum bias $\hat{B}$ (log scale) for each scatter estimator under different sample sizes $n$ and contamination levels $\varepsilon$ for dimension for dimension $p=15$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 1: Boxplots of the biases $\{\hat{b}_{k}^{(r)}\}_{1\leq r\leq R}$ (log-scale) through the $R=50$ replicates for each scatter estimator with $p=2$, $n=20$ and $k=1$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 2: Boxplots of the biases $\{\hat{b}_{k}^{(r)}\}_{1\leq r\leq R}$ (log-scale) through the $R=50$ replicates for each scatter estimator with $p=2$, $n=20$ and $k=20$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 3: Boxplots of the biases $\{\hat{b}_{k}^{(r)}\}_{1\leq r\leq R}$ (log-scale) through the $R=50$ replicates for each scatter estimator with $p=15$, $n=150$ and $k=0$
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 4: Boxplots of the biases $\{\hat{b}_{k}^{(r)}\}_{1\leq r\leq R}$ (log-scale) through the $R=50$ replicates for each scatter estimator with $p=15$, $n=150$ and $k=25$
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 5: $\hat{b}_{k}=\text{median}_{1\leq r\leq R}\{\hat{b}_{k}^{(r)}\}$ (log-scale) versus $k$ for each scatter estimator, under contamination levels $\varepsilon$. Dimension $p=2$ and $n=20$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 6: $\hat{b}_{k}=\text{median}_{1\leq r\leq R}\{\hat{b}_{k}^{(r)}\}$ (log-scale) versus $k$ for each scatter estimator, under contamination levels $\varepsilon$. Dimension $p=2$ and $n=1000$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 7: $\hat{b}_{k}=\text{median}_{1\leq r\leq R}\{\hat{b}_{k}^{(r)}\}$ (log-scale) versus $k$ for each scatter estimator, under contamination levels $\varepsilon$. Dimension $p=10$ and $n=100$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 8: $\hat{b}_{k}=\text{median}_{1\leq r\leq R}\{\hat{b}_{k}^{(r)}\}$ (log-scale) versus $k$ for each scatter estimator, under contamination levels $\varepsilon$. Dimension $p=10$ and $n=5000$.